OpenStudy (anonymous):

subtract and simplify √18-√8

OpenStudy (anonymous):

\[9\sqrt{2}-2\sqrt{2}\]=\[7\sqrt{2}\]

OpenStudy (campbell_st):

\[3\sqrt{2} -2\sqrt{?}\]

OpenStudy (anonymous):

yeah @campbell_st is right i forgot to take the root of the 9

OpenStudy (anonymous):

\[5\sqrt{2} \] would it be that or just what you put?

OpenStudy (anonymous):

it would just be \[\sqrt{2}\]

OpenStudy (anonymous):

oh true nevermind i forgot it was subtracting